OpenStudy (anonymous):

A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (anonymous):

can do it!...use x=1,2,3,4,sustitute in the equation and find the y, draw the graph...and then find x when y = 550...do u wnt me to help u further

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Note there is an error in the question...its not y=35.3x, it is y=35 . 3^x y= 35 . 3^x (35 is the initial number of bacteria) y= 35 . 3^(1) (substitute when x=1, so that u can graph it) y=105 when x=1 (these r the points which u r goin to plot in the graph) then y=35.3^x (substite when x=2) y = 35 * 3^2 y = 35 * 9 = 315 then do the same thing and u get a graph of x/y axis, plot the points and extend the curve and then draw a line(dotted) from the y=550 and chck x axis= which is the hours.... (sorry for the delay, cuz ur question ws wrng)

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

np:)

OpenStudy (anonymous):

thats not right..i used that and it marked my answer wrong